dy/dx=e^(x+y) -1.Please solve this differential equation.

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
Put x+y=z
So dy/dx+1= dz/dx
Or dy/dx= dz/dx-1
So the original equation becomes
dz/dx-1= e^z-1
Or x+e^-z= C
Or x+e^-(x+y)= C